# Resolving the Exploitation-Exploration Dilemma in Evolutionary Algorithms: A Novel Human-Centered Framework

### v1

@@ 5.1 Discussion
Throughout these experiments,various configurations of the HSSCP were employed during the local search phase, with each configuration corresponding to a distinct sequence-generative rule. For clarity, details of experiments involving sequence-generative rules that yielded unsatisfactory results in comparison to the global search phase are omitted. Instead, this section focuses on configurations that proved effective for each problem instance. The reason is, unsuccessful attempts in heuristics cannot provide us with information according to which one can infer any result general enough about the function in question or sufficient to theorization. The results of the global search phase are presented in table (1).
For the local search phase, the HSSCP configuration involved the following steps:
@@
As demonstrated in Tables (1, 3, 3, 8, 8, 8, 8, 8), the integration of the GA within the HCTPS framework consistently outperformed the standalone GA in terms of best objective values and statistical measures. Through the HCTPS-GA, we managed to find solutions that are within a tight vicinity of the known global optimum for each objective function, and even achieved the exact global optimum for F11 and F12.
Assessment of the HCTPS framework
The HCTPS is evaluated against the traditional approach based on the following key criteria: achieving the goal of maximizing exploration without minimizing exploitation,and the approach of achievement, and the invested cost vs produced value.
@@
    Goal achievement, and approach: The HCTPS framework effectively resolves the inherent conflict between exploration and exploitation that is characteristic of the traditional framework. This resolution allows for maximized exploration without compromising exploitation which is fundamental to the HCTPS.
@@
    Cost-benefit analysis: The HCTPS focuses tuning efforts towards the HSSCP, which is the core mechanism of this framework. While the tuning process may require trial and error to identify promising subcubes in the search space, the benefits significantly outweigh the tuning efforts cost. Not only that through the HSSCP, the limitations of the traditional approach listed in section (3), are effectively circumvented, it is the only reliable way to circumvent them. Also it is a more informed decision to invest more tuning efforts in configuring the HSSCP rather than internal EA parameters given the unintelligible interdependencies between internal parameters of the algorithm and the ambiguity of their specific roles in driving exploration versus exploitation internally. Even when specific HSSCP configurations fail to yield satisfactory results to the user, they still provide valuable insights into the search space by identifying regions that may not need further exploration. Furthermore, the intuitive and straightforward nature of fine-tuning the HSSCP enhances its practicality. This makes the HCTPS framework both an effective and efficient alternative to the traditional one for achieving the desired balance between exploration and exploitation.
6 Conclusions
The HCTPS framework represents a significant advancement in addressing the limitations of the traditional approach to managing the exploitation-exploration dynamics in EAs. By resolving the inherent conflict between these two goals, the framework is thus able to maximize exploration without diminishing exploitation, a necessary condition for effective search navigation.
@@ 6 Conclusions
While the fine-tuning of the HSSCP may involve an element of trial and error, this process is neither unique to HCTPS nor a hindrance to its adoption. EAs universally require trial and error experimentation to tailor parameter values to specific problem contexts. The HSSCP, however, offers much more valuable returns regarding maximizing exploration without introducing unnecessary randomization to the structure of the algorithm, in addition to its adaptability and intuitive implementation. Moreover, the ability to parallelize sub-cube testing provides a practical avenue for alleviating the computational burden, enabling concurrent exploration of multiple regions and expediting the search process.
Furthermore, the HCTPS framework applicability is not limited to EAs; it can be applied to any search algorithm as balancing exploration and exploitation is the fundamental principle to effective search.
@@ 7 Future Work
Moving forward, promising avenues for research could potentially lie in advancing human-centered tuning techniques within the HCTPS framework. Specifically, efforts can focus on developing generalizable rules for partitioning the search space to enhance exploration potential, particularly in high-dimensional contexts.
@@ Acknowledgments
I sincerely thank the organizers of the SIAM-MDS24 meeting for providing me with the opportunity to present and discuss a poster on my research, which later evolved into this paper.

### v2

@@ 6.1 Discussion
Throughout these experiments, various configurations of the SSCP were employed during the local search phase, with each configuration corresponding to a distinct sequence-generative rule. For clarity, details of experiments involving sequence-generative rules that yielded unsatisfactory results in comparison to the global search phase are omitted. Instead, this section focuses on configurations that proved effective for each problem instance. The reason is, unsuccessful attempts in heuristics cannot provide us with information according to which one can infer any result general enough about the function in question or sufficient to theorization. The results of the global search phase are presented in table (1).
For the local search phase, the successful SSCP configuration involved the following steps:
@@
As seen in Tables (1, 3, 3, 8, 8, 8, 8, 8), the results of the integration of the GA within the HCTPS framework outperforms the standalone GA in terms of best objective values and statistical measures. Through the HCTPS-GA, we managed to find solutions that are within a tight vicinity of the known global optimum for each objective function, and even achieved the exact global optimum for F11 and F12. However, it is important to emphasize that the success of reaching a tight neighborhood of the global optimum or finding the global optimum itself for a specific problem within the HCTPS depends on finding the appropriate sub-cube $J$ for the search algorithm $g$ in use, and since $J$ cannot be known a priori especially when dealing with black box optimization problems, the user may need to experiment with multiple sequence-generative rules before identifying an optimal SSCP configuration in favor of finding the right sub-cube.
7 HCTPS Assessment
We evaluate HCTPS in light of the criteria that define a systematic resolution to the exploration-exploitation dilemma in the introduction section (1), namely:
@@ 7 HCTPS Assessment
    Decoupled and independent control of exploration and exploitation
@@
    Providing explicit and adaptive regulation of exploration and exploitation
  3. 3.
    Theoretical and computational feasibility
In addition, we also consider the cost–benefit implications of adopting HCTPS in practice. Below, we summarize how HCTPS addresses each point and examine its comparative advantages over the current approach.
7.1 Decoupled and Independent Control
A principal contribution of HCTPS is that it decouples exploration from exploitation by shifting the regulatory burden to an external parameter—the SSCP. Concretely:
  • Global Phase: The EA’s inherent exploitation mechanisms (e.g., its population convergence tendencies, mutation/selection synergy, etc.) remain intact and need not be modified or “loosened” to increase exploration. Instead, the algorithm simply executes as usual across the entire search cube
  • Local Phase: Exploration is selectively and additively intensified in chosen sub-cubes $\{I_{i}\}_{i=1}^{n}$ of $I$. By letting the user dynamically specify or adjust which sub-cubes to explore and to what extent, HCTPS ensures that raising exploration does not implicitly undermine the EA’s convergence properties.
This external and human-centered mechanism fundamentally differs from the traditional approach (i.e., re-tuning internal operator parameters). Exploration and exploitation thus become governed by distinct, independently adjustable mechanisms: the EA’s built-in exploitation potential, on the one hand, and the SSCP-based search-space relocation, on the other.
7.2 Explicit and Adaptive Regulation
An equally crucial requirement for a systematic resolution is that exploration be regulated explicitly and be adaptive to real-time observations. HCTPS satisfies these in the following ways:
  1. 1.
    Explicit parameter (SSCP): The SSCP is the dominant and only parameter the user needs to focus on to modulate exploration. Rather than distributing tuning efforts among mutation rates, crossover probabilities, or other interdependent parameters, the user exerts direct control over sub-cube selection and size (including possible scaling factors).
  2. 2.
    Adaptive local search: The user can progressively generate additional sub-cubes $\{I_{i}\}$ or refine existing ones based on observed partial results (e.g., fitness improvement rates, solution clustering, .. etc). This selective local-phase configuration (Step 2 and Step 3 in procedure (1) thus adapts the exploration intensity to real-time insights.
By centralizing exploration decisions into one external “knob,” HCTPS avoids entangled parameter interdependencies, thereby enabling a principled approach to increasing or reducing exploration as the search unfolds.
7.3 Theoretical and computational feasibility
HCTPS meets the third requirement for a systematic resolution by remaining algorithm-agnostic, not imposing exponential expansions in representation size, nor mandating extensive operator re-tuning:
  • Algorithm-agnostic: HCTPS is a procedural add-on that can be combined with any evolutionary or non-evolutionary search algorithm $g$. The only assumption is that $g$ can run on a given subregion of the search space.
  • No exponential increase in representation set: Unlike attempts to boost exploration by lengthening encodings or using prohibitively large populations, HCTPS redefines which portion of $I$, the algorithm sees at each local phase, avoiding unbounded expansions in representation cardinality.
  • No exhaustive parameter sweeps: By focusing all exploration tuning on the SSCP, HCTPS circumvents the need to balance multiple, unintelligibly interdependent parameters. This single-parameter approach is less prone to combinatorial parameter inflation.
Moreover, the theoretical guarantee in Theorem 1 establishes that HCTPS strictly improves exploration coverage over the traditional approach, provided the chosen sub-cubes in the local phase are sufficiently diverse.
7.4 Cost–Benefit Analysis
Although the user (i.e., decision-maker) invests time in constructing or refining an SSCP-based sequence of sub-cubes, this investment replaces the extensive trial-and-error that typically arises when tuning many internal EA parameters simultaneously:
  1. 1.
    Focused tuning effort: All exploration-specific customizations are consolidated into one parameter class (SSCP). This spares the user from contending with unintelligible interdependencies among operators, probabilities, and population dynamics.
  2. 2.
    Incremental refinement: Even when certain SSCP configurations fail, they often yield search-space insights regarding which subregions are not promising. This is arguably more interpretable than concluding that “some parameter triple $(N,p_{c},p_{m})$ was suboptimal” in a black-box sense.
  3. 3.
    Parallelizability: Since each sub-cube can be handled in an isolated local phase, the user may run local-phase searches in parallel if computational resources permit, further reducing overhead.
Ultimately, the gains in exploration efficacy and interpretability of the local-phase results typically outweigh the cost of specifying or updating SSCP rules.
8 Summary, and Key Takeaways
8.1 Summary
The fundamental obstacle identified in the current approach practiced in the literature in dealing with the exploration-exploitation dilemma arises from two interacting issues:
  1. 1.
    The fixed representation set $X$ size in which the EA operates
  2. 2.
    The unintelligible interdependencies between its traditional parameters
(1) alone implies that, at best, the algorithm will only be able to unlock a subset from $I$ with the same cardinality of $X$, and whose cardinality is a function of the pre-determined parameter $L$ (solutions bit-encoding length). While (2) alone implies that the collective contribution of the internal mechanics of the EA is not predictable. (1) and (2) together makes it very reasonable to overlook potential promising regions or solutions if they happen to not exit in the subset we are to unlock. Increasing exploration under (1) and (2) in the current approach necessitates increasing the size of $X$ and population size $N$ as perminilaries and the consequences is a huge computational burden.
To avoid such consequences, we proposed a two-phase search framework to which the human-user serves as a meta-paramter, thus the name ”human-centered two-phase search” (HCTPS). This framework allows for the free adaptive adjustment of exploration without affecting the inherent exploitation of the algorithm in use, thus make it possible to reach the proper exploration to exploitation ratio that fit the specific problem under consideration and the underlying algorithm in use, and while being theoretically and computationally feasible, and thus resolves the exploration exploitation dilemma.
In section (5), we detailed the search process within the proposed framework, which restructures the search into two consecutive phases: a global search phase followed by a local search phase. In the first, the EA operates over the entire search cube $I$ for $k$ iterations, and in this the EA distributes its inherent exploration-exploitation potential over $I$. In the later, the search is carried out over selected partitions ”sub-cubes” of $I$ each for the same $k$ iterations, and in this the EA distributes its inherent exploitation-explortion potential over different-in-size sub-regions of $I$, and the latter phase continues until satisfactory results are identified. The rigorous implementation of the local search phase is facilitated through the introduction of SSCP which follows a human-centered tuning strategy (see section 2.2), and a compact description of the HCTPS was provided by procedure (1).
8.1.1 HCTPS Effect on EAs as Global Optimizers
Resolving the exploration-exploitation dilemma in EAs is a prerequisite for enhancing their utility as global optimization tools—the latter cannot follow without the first. Therefore, applying an arbitrary EA (i.e., any EA with some fixed parameter values) within the HCTPS and comparing against the same EA as standalone serves as a nice way to illustrate its concrete implementation, and demonstrate its effectiveness against the current approach (i.e, focusing all tunning efforts on the SSCP vs distribution over traditional EA parameters), and this was carried out in section (6) with the canonical GA as an algorithm choice.
The aim of the experiment is not to serve as a tie-breaker to the framework effectiveness which means an unsuccessful SSCP configuration doesn’t imply that the framework is defective but rather that the specific SSCP configuration is not the right one for the specific algorithm in use and the problem it is applied to. The framework’s guaranteed resolution of the dilemma in question, and thereby increasing the chances of hitting a tight approximation of global optima is supported by its theoretical and conceptual foundations, as presented in section (5.3).
8.2 Key Takeaways
The key takeaways from this work can be summarized as follows:
  1. 1.
    Reconciling exploration–exploitation: Within the HCTPS framework, exploration and exploitation are not inherently conflicting because they are decoupled. This allows for the maximization of the first without diminishing the latter, a necessary condition for effective search performance.
  2. 2.
    Significantly enhanced coverage: Embedding any EA within HCTPS significantly increases search coverage and outperforms the same EA without HCTPS. The theoretical foundation shows that decoupling exploration in this way strictly expands the set of points the algorithm can effectively sample, thus reducing the chances of missing promising niches.
  3. 3.
    Single-parameter focus: A key advantage of HCTPS is that all exploration-related tuning efforts are directed towards a single parameter (the SSCP), thereby removing many of the unintelligible, compounded interactions among traditional EA parameters.
  4. 4.
@@ 8.2 Key Takeaways
    Adaptability and trial-and-error: While the fine-tuning of the SSCP may involve an element of trial and error, this is neither unique to HCTPS nor a hindrance to its adoption. EAs universally require trial and error experimentation to tailor parameter values to specific problem contexts. The human-centered SSCP, however, offers much more valuable returns regarding maximizing exploration without introducing unnecessary randomization to the structure of the algorithm in use in addition to its adaptability, simple tuning process and intuitive implementation. Moreover, the ability to parallelize sub-cube testing provides a practical avenue for alleviating the computational burden, enabling concurrent exploration of multiple regions and expediting the search process.
  5. 5.
    Broad applicability: The HCTPS framework applicability is not limited to EAs; the general idea of “decouple and externally direct exploration” applies to any search method, not just EAs. Whenever an algorithm’s default parameters alone fail to explore sufficiently, an external “search-space control” layer can fill that gap.
  6. 6.
    In closing, the HCTPS provides an overarching reconcilatory framework that unifies exploration and exploitation without compromising either one. By disentangling these two processes—and placing exploration under dynamic human-centered control—the framework has the potential to offer a robust, extensible means of improving global search outcomes across a wide spectrum of search computational problems prone to the exploration exploitation dilemma.
@@ 9 Future Work
Future research could potentially focus on developing interactive systems where humans and automated agents work in collaboration to refine sequence-generative rules for search space partitioning. One promising direction involves designing user interfaces that provide real-time visualization of the search space and the current partitioning strategy. Such interfaces would display performance metrics and spatial distributions of explored subcubes, enabling users to quickly identify regions of interest or redundancy.
@@ Acknowledgments
I sincerely thank the organizers SIAM-MDS24 for the opportunity to present and discuss early phases of this work, professor Rana Dajani (Hashemite University) for clarifying evolutionary biology concepts which helped me refine some paragraphs in the introduction section, and professor Thomas Bäck (Leiden University) for his comments on the early version of this manuscript, which inspired improving clarity in the presentation of its ideas in this version.